OpenStudy (anonymous):

What is the remainder of (x^3 – 6x – 9x + 3) ÷ (x - 3) A. -51 B. -51/x - 3 C. -17/x D. -17/ x - 3

OpenStudy (anonymous):

someone please help me??????

OpenStudy (helder_edwin):

\[ 3^3-6\cdot 3-9\cdot 3+3= \]

OpenStudy (helder_edwin):

isn't there missing an x^2?